OpenStudy (anonymous):

Help please.How do you decide which trigonometric ratio to use? How do you know which side is the opposite, adjacent, or hypotenuse based on the given information in each problem? Be specific in your explanation.

Nnesha (nnesha):



Nnesha (nnesha):

hypotenuse is the longest leg of right triangle adjacent and opposite sides depends on the angle you are looking at adjacent: the side the touche the angle (theta)

OpenStudy (jhannybean):

SOHCAHTOA \[\sf sin (\theta) = \frac{opp}{hyp} = SOH \]\[\sf cos(\theta) = \frac{adj} {hyp} = CAH\]\[\sf tan(\theta) = \frac{opp}{adj} = TOA\]

Nnesha (nnesha):

like for this angle x  \[\rm sin \rm \theta = \frac{ opposite }{ hypotenuse }~~~~ \cos \theta = \frac{ adjacent }{ hypotenuse } ~~\tan \theta = \frac{ opposite }{ adjacent }\]